How do you find three solutions to the equation y=8x+5
Found 3 solutions by richwmiller, solver91311, stanbon:
Answer by richwmiller(17219)

the easiest is when x=0 then y=5
when x=1 then y=13
when x=2 y=21

Answer by solver91311(24713)



Pick a value for . It can be anything you like, but small integers generally make for simpler arithmetic and less chance of error. Substitute the value you chose in place of in the equation. Do the arithmetic to determine the value of that results. Create an ordered pair, using the value of you chose and the value of you calculated. That is one solution.

Repeat two more times using different values for each time.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

How do you find three solutions to the equation y=8x+5
-------
Let x = 0, then y = 8*0+5 = 5
------
Let x = 2, then y = 8*2+5 = 21
-----
Let x = -3, then y = 8*-3+5 = -19
=====================================
Cheers,
Stan H.
====================